A capacitor consists of two closely spaced metal conductors of large area, separated by a thin insulating foil. It has an electrical capacity of 3400.0 μF and is charged to a potential difference of 81.0 V. Calculate the amount of energy stored in the capacitor. Submit Answer Tries 0/99 Calculate the charge on this capacitor when the electrical energy stored in the capacitor is 12.00 J. Submit Answer Tries 0/99 If the two plates of the capacitor have their separation increased by a factor of 6 while the charge on the plates remains constant, by what factor is the energy stored in the capacitor increased? Submit Answer Tries 0/99
Answer #1

answer) we know the formula

a)E=1/2cv2=1/2*3400*10-6*812=11.15J

so the answer is 11 J or 11.2 J or 11.15 J

b) we know the formula

E=Q2/2C=12 J

Q2=2*3400*10-6*12=0.0816

Q=0.286 C

so the answer is 0.286 C or 0.29 C or 0.3C

c) we have the formula

E=Q2/2C

C=\varepsilonA/d

when the plate separation increases by a factor of 6 then capacitance decreases by a factor of 6 and since capacitance is inversely proportional to energy then the energy increases by factor of 6

so the answer is 6.